# Assessing AI vs Human-Authored Spear Phishing SMS Attacks: An Empirical Study Using the TRAPD Method

### v1

@@ 4.2.2 AI Generation
A script was written to call the GPT-4 API to generate spear phishing message outputs automatically. The same three prompts given to the student authors for each target were fed to the script to generate the AI-generated spear phishing messages. An example of a GPT-4 output is shown below:
@@
Since GPT-4 outputs are not completely reproducible, three different responses were gathered from each prompt. This resulted in 9 messages generated by the model for each target. Similar to what was done with the human-authored messages, we selected at random 2 messages for each topic, totaling 246 AI-generated messages that will be used for the next phase.
After both sub-phases completed, each target had 12 spear phishing messages, 6 made from GPT-4 and 6 were human-made. By the end of this phase 492 simulated spear phishing “attacks” were prepared for assessment.
(a) Rank Distribution of AI vs Human Messages
(b) Probability distribution of AI performance
Figure 2: AI vs Human performance results (Rank & Click Probability.
4.3 Target Interview
Of the 41 targets that signed up to participate, 25 were able to voluntarily come for the target interview. In this interview, targets were shown the personalized spear phishing messages from the message generation phase. At this point, they were not informed that some messages were AI-generated. Each participant consented to be recorded for future analysis. We break down this phase according to the steps in the TRAPD methodology:
@@ 4.3.1 Threshold rank order
The 12 targeted messages were shown to each target during the interview, with all messages displayed on a separate printed piece of paper. Participants were asked to arrange messages, ranking them on their ability to convince them to click on the link provided. Figure 1a shows an example of how the targets laid out the messages in the experiment, illustrating the ranked order from most to least convincing, as well as the threshold where they would have started to click the link based on the participants’ assessments.
@@ 4.3.2 Qualitative Assessment
Participants were then asked what elements in the messages led them to be tricked, as well as other feedback that they were willing to share regarding the message. They were also asked what made them more cautious of the less convincing messages.
@@ 4.3.3 Label as Human vs AI
Participants were then notified that one or more of the messages were created by an AI, and were asked to identify which messages they believe to be human-made vs computer generated. Figure 1a shows an example on how the targets accomplished this, by placing AI “markers” on the messages that they believed to be AI-generated. Similar to the previous part of the interview, they were then asked what made them think the messages they chose were made by humans or AI. At the target’s request, researchers revealed which messages were made by AI and human authors.
Audio recordings were made for all 25 interviews which were then assimilated for transcription and further analysis.
@@ 4.4 Statistical Methodology
Several statistical analyses were conducted to evaluate the difference in performance of AI-generated and human-authored spear phishing messages (RQ1), the difference in performance of particular elements in the spear phishing messages (RQ2), the overall accuracy of the targets in identifying AI-generated messages (RQ3), and whether certain factors in the messages helped them identify AI-generated messages better (RQ4).
To address RQ1, we began by comparing the average ranking of both types of messages across all the targets. To verify statistical significance, we utilized a permutation test to validate our findings. Following this, we applied a logistic regression model to predict the likelihood of a subject clicking on a link based on whether the message was AI or human. To account for potential correlations within subjects and authors, we fitted a generalized linear mixed-effects model (GLMM). To assess how similar or better the performance of AI messages are compared to human messages, we also employed a Bayesian logistic model, using a horseshoe prior on the log odds ratio, pulling estimates towards zero when evidence was weak but not excessively when evidence was strong.
To address RQ2, we compared the average rankings of the messages according to each of the three topics: Job, Hobby or Social. We also used a logistic regression model in relation to the three topics to assess whether a topic was more successful in generating clicks than the others.
@@ 5.1 Statistical Analysis
In this section we report on the statistical results as from the models described in the methodology. The statistical models and methods used were derived from data gathered during the target recruitment and interview phases.
@@
To assess performance between AI and human messages (RQ1), we compared their performance based on how high they were ranked in the target interviews and how often each type of messages would be clicked by out targets. On average, AI-generated messages ranked slightly higher (6.41) than human-authored messages (6.58). Figure 2a describes the distribution of the AI and human messages across each rank, as well as their averages.
Results from our statistical analysis indicated no significant difference between the two groups, with a p-value of 0.665. To account for the non-normal distribution of ranks and the variation across targets, a permutation test was conducted, but it again showed no significant difference in ranks between AI and human-generated messages.
@@
We employed a Bayesian logistic model to further the analysis. Using a horseshoe prior on the log odds ratio, we obtained a posterior mean odds ratio of 1.22 (95% CI: 0.86 to 2.05) with a median posterior of 1.16. From this posterior distribution, we estimated that the probability of the odds ratio being less than 1 (indicating humans are better than AI) is 19.7%, while the probability of it being greater than 1 (indicating AI is better than humans) is 80.3%. The probability of the odds ratio exceeding 1.5 (AI being more than 50% better than humans at generating clicks) is only 19.9%. A plot of the posterior distribution of the odds ratio shown in Figure 2b indicates that, while there is a long right tail suggesting a possibility of AI being significantly better, most of the probability mass is near 1, suggesting neither being definitively better.
Figure 3: Rank Distribution of Each Topic
@@ 5.1.2 Topic-based Ranking & Click Probability
One of the key content characteristics of the messages (RQ2) relates to their topic: Job, Hobby, or Social media post. On average, Job ranked the highest (5.71), followed by Hobby (6.66) with Social ranking the lowest (7.13). The distribution of these rankings is shown on Figure 3. Using a Tukey post-hoc test to analyze statistical significance of the ranking of each topic, we derived the confidence intervals of each pair of topics. Results show that the mean rank for Job is significantly higher than Social, while Job-Hobby is not statistically significant, and much less for Hobby-Social.
@@
TABLE II: AI Identification Confusion Matrix
@@
AI 78 72 150
Human 72 78 150
SUM 150 (52%) 150 (52%) 156/300 (52%)
5.1.3 Identifying Message Origin
To assess if subjects could correctly identify whether a message was AI-generated or human-authored (RQ3), we analyzed their guesses for both types of messages. Subjects correctly identified the message origin 52% of the time. Note that 50% would be expected from randomly guessing. Table II describes the confusion matrix that details their accuracy.
@@ 5.2 Content Characteristics of Persuasive Messages
This section addresses RQ2 from a qualitative perspective. Specifically, the following subsections identify the core themes that resulted from our analysis of the verbal explanations participants gave when describing why they sorted messages as most likely to deceive them or least likely to deceive them.
@@ 5.2.1 URL
The URL was mentioned by 64% of all participants as having convincing attributes (16/25). Six participants noted how the URL can provide a sense of trust, with one noting, “BYU, I kind of trust BYU and the URL here” (T41). Three participants mentioned convincing characteristics with regards to the domain of the URL. One stated, “The biggest part of this one [URL], and I even debated whether I would put it first or second; but it has an .edu link . . . and maybe I’m very ignorant about this, but I feel like that’s harder to, like, create a fake .edu website” (T16). In contrast, another participant used similar reasoning for not clicking on a link, noting, “At first, I thought, Oh, I would click on that. But then it says google.com.org; which is weird, because I’ve never seen that before” (T07). This same participant mentioned some inaccuracies related to the domain that would cause him/her to rethink clicking on it. Another participant mentioned the use of a .net website which contributed to them clicking on the link because the domain seemed trustworthy.
Two participants were more likely to click on the link if it included the HTTPS protocol. One said that “If it looks legit, yeah, like HTTPS and the /ProvoLibrary, something like that” (T41), then they would click. Another participant said, “they all say HTTPS, which makes me feel like it’s secure” (T07). Personal association with a URL may also play a role, with one participant stating s/he was more likely to click on the link if it included his name within the URL saying that “the link actually has my name in it. That would have really like, definitely thrown me for a loop, like really made me like, actually think about clicking the link just because it’s like, oh, this is actually from maybe from BYU” (T20).
@@ 5.2.2 Technology Communication Medium
The technology medium used for communication was mentioned by 40% (10/25) of the interviewees. One of the 10 mentioned the medium increasing their likelihood of clicking on the link. S/he mentioned that it is “not that abnormal for [them] to receive” text messages from their home city” (T13). However, most targets saw the text medium as a red flag, even when they were inclined to believe the content. This was sometimes because of warnings from trainings they had received. One explained that “at work we were warned that all, like, valid messages will be through emails” (T25) and “This one, [employer] is not going to reach out to me via text for security stuff, period” (T08). T37 mentioned that they don’t have a “company texting deal,” making them more suspicious of SMS messages. Another men-tioned that “it’s very rare that people message me, you know, we talk over Slack or we talk over WhatsApp, maybe through email” (T31). Others noted that other communication tools, such as an online learning platform (T09) or a genealogy website (T02), would not have sent them a text message. Interestingly, in all these cases, the technology communications medium was associated with the source of the message. In other words, receiving a text message was not abnormal in and of itself, but receiving a text message from a particular source was deemed either appropriate or not.
5.2.3 Context Inaccuracies
The presence of inaccuracies in the messages was mentioned by 28% (7 of 25) of the targets. These inaccuracies contributed to perceptions about message credibility. Participants identified discrepancies related to the names or characteristics of individuals or enti-ties within their professional spheres. For example, two targets remarked about the message mentioning a col-league that does not exist, stating “there’s no Mike at work” (T37) or “there’s no one named Sarah on the BYU instructional design team” (T18). Another participant expressed confusion about being questioned on matters not relevant to their professional duties, saying that they “don’t deal with payables. So why is that asking me regarding payment?” (T25). Another target noted discrepencies related to job responsibilities, asserting that such library-related decisions would “go through me at the library”, making any logistical deviation “immediately sus-picious” (T01). Regarding messages related to their hobbies, one participant raised skepticism about the message claiming a “great deal for a hiking trail” which the participant knows through experience doesn’t actually require payment for use (T30). Another target was also suspicious of the message being from a library that they “just don’t use… so they would not know me to contact me” (T08). In summary, the instances where inaccuracies were mentioned ranged from misrepresented personnel and responsibilities to factual discrepancies about the targets’ affiliations, activities, and expectations.
5.2.4 Personal Relevance
The messages being personally relevant (or irrelevant) to the participants was mentioned in 76% (19 of 25) of the interviews. Seventeen targets explained that personal relevance of the message made it more convincing. Several participants pointed to messages directly tied to their occupational responsibilities, with one saying that “because that is my job is to help people with records…this is one that I feel like I’m most likely to engage with” (T13). Another target, who is a banker at a credit union, noted the similarities between the targeted message and messages s/he receives at work. S/he noted, the “alert literally looks like the alert we get [at work] when there’s a fraud” (T33). Some participants pointed out how the spear phishing messages aligned with their personal interests. One expressed enthusiasm about a message describing a “Vineyard gardening club”, believing that “somebody from our community” may have sent it. They expressed that they “would love to get involved in that” (T31). While most participants were more convinced when messages were relevant to them, one more skeptical participant mentioned that although “this interests me and it could exist, but I’m not gonna click on it” (T08).
@@ 5.2.5 Scarcity Principle
The remaining participants (6 of 25) said they were less likely to click on the link if it included tactics such as fear and urgency. These participants explained that time sensitive or urgent messages feel like a warning sign. “That’s always a way of like, okay, they’re trying to get you to shut off your logical brain. Put yourself into ’Oh, my gosh, we have to do this right now.’ Yeah. Right.” Another participant confirmed, “If it seems urgent to you, you’re not going to click it because it’s very dangerous” (T13). Similarly, T41 explained, “I don’t like anything where it’s like hurry, offer and send because that always makes me warry” (T07). Similar to the their concerns with time-sensitive messages, participants also discussed how fear inciting messages can cause them to logically analyze the message at a deeper level. One explained, “To the message itself, I was a little bit unsure about just like, the fact that it was so time sensitive, and that it was like, kind of threatening to lock me out of all of my educational accounts if I didn’t do what it asked for. Because that’s not typically what I’ve seen be the case” (T16). And another stated, “Yeah. And it’s the idea of a new virus for young dogs. I’m going that’s fear mongering” (T01). The scarcity principle has long been used in literature to deceive readers. For the spear phishing messages in this study, participants had mixed reactions. Some felt the intended pressure to click, while others saw the language as manipulative and as a warning sign marking potential danger.
5.2.6 Messaging Style
A quarter of respondents (6 of 25) discussed the role of emojis in the context of message style. All six asserted that emojis diminished the credibility of the message. T34 expressed that “I think a lot of emojis is kind of something that usually is a red flag for me.” Multiple respondents conveyed that they did not anticipate receiving emoji-filled messages from senders claiming to represent professional organizations. For instance, T27 said that in the case of a legitimate organization such as a city biking club, emojis would be “out of place.” Similarly, T20 explained that if a message claimed to be from their university, the presence of emojis would cause them to feel “thrown off.” From the perspective of participants, the use of emojis in messages, particularly those seemingly sent by professional groups, caused such messages to lose legitimacy.

### v2

@@ 4.2.2 AI Generation
A script was written to call the GPT-4 API to generate spear phishing message outputs automatically. The same three prompts given to the student authors for each target were fed to the script to generate AI-generated spear phishing messages. An example of a GPT-4 output is shown below:
@@
Since GPT-4 outputs are not completely reproducible, three different responses were gathered from each prompt. This resulted in 9 messages generated by the model for each target. Similar to what was done with the human-authored messages, we selected at random 2 messages for each topic, totaling 246 AI-generated messages to be used for the next phase.
After both sub-phases were completed, each target had 12 spear phishing messages: 6 created by GPT-4 (2 for each of the 3 topics) and 6 created by humans (2 for each of the 3 topics). By the end of this phase, 492 simulated spear phishing “attacks” were prepared for assessment.
4.3 Target Interview and Sorting Activity
All 41 people who filled out the survey were invited to come back for the interview/sorting activity approximately a month later. Only 25 (61%) responded and arrived on their scheduled interview day. The distribution of participant demographics are shown on Figure 2. We were successful in recruiting a range of ages, as well as professions. Although 15 (60%) of the targets were affiliated with the university, only 8 (32%) were students. A range of jobs were reported (e.g., librarian, software engineer, instructional designer, sales agent, teaching assistant).
Figure 2: Target Demographics (n=25).
@@ 4.3.1 Threshold rank order
The 12 spear phishing SMS messages were shown to each target during the interview, with all messages displayed on a separate printed piece of paper showing a phone and text message image around the content. Participants were asked to arrange messages in order of most to least likely to lead to the target clicking on them. Figure 3 shows an example of how the targets laid out the messages in the experiment, illustrating the ranked order from most to least convincing, as well as the threshold where they would have started to click the link based on the participants’ assessments. Participants were not explicitly told the messages were spear phishing messages when they were asked to sort them.
Figure 3: A sample of the threshold rank ordering used in the target interview. After ranking the messages from most (red) to least (blue) likely, a marker (yellow) is placed to indicate the threshold for clicking.
@@ 4.3.2 Qualitative Assessment
Participants were then asked what elements of the messages led them to be tricked, as well as other feedback that they were willing to share regarding the criteria they used to sort messages. They were also asked what made them place the less convincing messages at the bottom. They were encouraged to look at individual messages when providing feedback in order to ground their comments with specific examples. This was often followed by more general assessments by targets about overall criteria they used that applied to more than a single message.
@@ 4.3.3 Label as Human vs AI
Participants were then told that "one or more of the messages were created by an AI," and were asked to place an AI "marker" on the message or messages they believe were created by AI as shown in Figure 4. They were then asked what made them think the messages they chose were made by humans or AI. Again, they were asked to use specific messages to ground their observations, followed by their overall thoughts. At the target’s request, researchers revealed which messages were made by AI vs human authors.
Audio recordings were made for all 25 interviews which were then assimilated for transcription and further analysis. Images of all sorting and labeling were also captured for data analysis.
Figure 4: An example on the AI labeling phase of the target interview.
@@ 4.4 Statistical Methodology
Several statistical analyses were conducted to evaluate the difference in performance of AI-generated and human-authored spear phishing messages (RQ1), the difference in performance of particular elements in the spear phishing messages (RQ2), the overall accuracy of the targets in identifying AI-generated messages (RQ3), and whether certain factors in the messages helped them identify AI-generated messages better (RQ4). All statistical analyses were carried out in R.
To address RQ1, we compared the average ranking of both types of messages across all the targets. To verify statistical significance, we utilized a permutation test [24] to validate our findings. Following this, we applied a logistic regression model [25] to predict the likelihood of a subject saying they would click on a link (i.e., being above the click threshold) based on whether the message was AI or human. To account for potential correlations within subjects and authors, we fitted a generalized linear mixed-effects model (GLMM).[26] To assess how similar or better the performance of AI messages are compared to human messages, we also employed a Bayesian logistic model[27], using a horseshoe prior on the log odds ratio, pulling estimates towards zero when evidence was weak but not excessively when evidence was strong.
To address RQ2, we compared the average rankings of the messages according to each of the three topics: Job, Hobby or Social. We also used a logistic regression model based on the threshold self-assessments in relation to the three topics to assess whether a topic was more successful in generating clicks than the others.
@@ 5.1 Statistical Analysis
The main metrics for performance of each topic (job, hobby and social) and source (Human and AI) is their average ranking as determined by the order which their corresponding messages were placed by the targets, and the click probability, as determined by the threshold marker. Table 1 shows a summary of the average ranking and click probability of each topic and source. This subsection expounds on the statistical significance of each result.
Topic Ave. Rank Click Rate
Job 5.71 38%
Hobby 6.66 19%
Social 7.13 17%
Source
Human 6.58 21.33%
AI 6.41 28%
Table 1: Average Rank and Click Rate by Topic and Source. Job-related and AI-generated messages generated more clicks than their counterparts.
@@
To assess performance between AI and human messages (RQ1), we compared their performance based on how high they were ranked in the target interviews and how often each type of message would be clicked by our targets. On average, AI-generated messages ranked slightly higher (6.41) than human-authored messages (6.58). Figure 5 describes the distribution of the AI and human messages across each rank.
Figure 5: Rank Distribution from each source. On average, AI-generated messages ranked slightly higher than human-generated messages.
Results from our statistical analysis did not indicate significant differences between the two groups, with a p-value of 0.665. To account for the non-normal distribution of ranks and the variation across targets, a permutation test was conducted, but it again showed no significant difference in ranks between AI and human-generated messages.
@@
We employed a Bayesian logistic model to further the analysis. Using a horseshoe prior on the log odds ratio, we obtained a posterior mean odds ratio of 1.22 (95% CI: 0.86 to 2.05) with a median posterior of 1.16. From this posterior distribution, we estimated that the probability of the odds ratio being less than 1 (indicating humans are better than AI) is 19.7%, while the probability of it being greater than 1 (indicating AI is better than humans) is 80.3%. The probability of the odds ratio exceeding 1.5 (AI being more than 50% better than humans at generating clicks) is 19.9%. A plot of the posterior distribution of the odds ratio shown in Figure 6 indicates that, while there is a long right tail suggesting a possibility of AI being significantly better, most of the probability mass is near 1, suggesting neither being definitively better.
Figure 6: Posterior distribution of the odds ratio of Human vs AI messages. Results suggest that AI-generated messages performed as well if not better than human-authored messages.
@@ 5.1.2 Topic-based Ranking & Click Probability
One of the key content characteristics of the messages (RQ2) relates to their topic: Job, Hobby, or Social media post. On average, Job ranked the highest (5.71), followed by Hobby (6.66) with Social ranking the lowest (7.13). The distribution of these rankings is shown in Figure 7. Using a Tukey post-hoc test to analyze statistical significance of the ranking of each topic, we derived the confidence intervals of each pair of topics. Results show that the mean rank for Job is significantly higher than Social, while Job-Hobby is not statistically significant, and much less for Hobby-Social.
Figure 7: Rank Distribution of Each Topic. Job-related messages ranked higher than hobby and social-related messages.
@@
5.1.3 Identifying Message Origin
To assess if subjects could correctly identify whether a message was AI-generated or human-authored (RQ3), we analyzed their guesses for both types of messages. Subjects correctly identified the message origin 52% of the time. Note that 50% would be expected from randomly guessing. Table 2 describes the confusion matrix that details their accuracy.
@@ 5.1.3 Identifying Message Origin
AI (truth) 78 72 150
Human (truth) 72 78 150
SUM 150 (52%) 150 (52%) 156 (52%)
Table 2: AI Identification Confusion Matrix. Results suggest that target accuracy is not much better than flipping a coin.
@@
5.1.4 Demographics and Phishing Experience
Participants in this study were roughly half male and half female and had various occupational roles including entry-level employees, mid-level managers, and college students. Despite this diversity, no statistically significant differences were observed across demographic groups for click rates or the ability to identify message origin (AI vs. human). Prior experience with phishing did not appear to significantly improve performance in identifying AI-generated messages or reducing click rates, aligning with prior research suggesting that familiarity alone may not translate to better detection.[28]
@@ 5.2 Content Characteristics of Persuasive Messages
This section addresses RQ2 from a qualitative perspective. Specifically, the following subsections identify the core themes that resulted from our analysis of the verbal explanations participants gave when describing why they sorted messages as most likely to deceive them or least likely to deceive them. Table 3 provides a summary of the percentage of targets who mentioned these themes that influence their perception of a convincing message.
Themes % Mentioned
Relevance 76%
URL 64%
Medium 40%
Style 40%
Scarcity 32%
Inaccuracies 28%
Rewards 28%
Sender 28%
Table 3: Key Content Characteristics Influencing Message Persuasiveness. Most messages that were convincing to targets were identified to be "relevant" to them to some degree.
5.2.1 Personal Relevance
The messages being personally relevant (or irrelevant) to the participants was mentioned in 76% (19 of 25) of the interviews. Seventeen targets explained that personal relevance of the message made it more convincing. Several participants pointed to messages directly tied to their occupational responsibilities, with one saying that “because that is my job is to help people with records…this is one that I feel like I’m most likely to engage with" (T13). Another target, who is a banker at a credit union, noted the similarities between the targeted message and messages s/he receives at work. S/he noted, the “alert literally looks like the alert we get [at work] when there’s a fraud” (T33). Some participants pointed out how the spear phishing messages aligned with their personal interests. One expressed enthusiasm about a message describing a “Vineyard gardening club”, believing that “somebody from our community” may have sent it. They expressed that they “would love to get involved in that" (T31). While most participants were more convinced when messages were relevant to them, one more skeptical participant mentioned that although “this interests me and it could exist, but I’m not gonna click on it” (T08).
@@ 5.2.2 URL
The URL was mentioned by 64% of all participants as having convincing attributes (16/25). Six participants noted how the URL can provide a sense of trust, with one noting, “[my university], I kind of trust [my university] and the URL here" (T41). Three participants mentioned convincing characteristics with regards to the domain of the URL. One stated, “The biggest part of this one [URL], and I even debated whether I would put it first or second; but it has an .edu link . . . and maybe I’m very ignorant about this, but I feel like that’s harder to, like, create a fake .edu website” (T16). In contrast, another participant used similar reasoning for not clicking on a link, noting, “At first, I thought, Oh, I would click on that. But then it says google.com.org; which is weird, because I’ve never seen that before” (T07). This same participant mentioned some inaccuracies related to the domain that would cause him/her to rethink clicking on it. Another participant mentioned the use of a .net website which contributed to them clicking on the link because the domain seemed trustworthy.
Two participants were more likely to click on the link if it included the HTTPS protocol. One said that “If it looks legit, yeah, like HTTPS and the /ProvoLibrary, something like that” (T41), then they would click. Another participant said, “they all say HTTPS, which makes me feel like it’s secure” (T07). Personal association with a URL may also play a role, with one participant stating s/he was more likely to click on the link if it included his name within the URL saying that “the link actually has my name in it. That would have really like, definitely thrown me for a loop, like really made me like, actually think about clicking the link just because it’s like, oh, this is actually from maybe from [my university]” (T20).
@@ 5.2.3 Technology Communication Medium
The technology medium used for communication was mentioned by 40% (10/25) of the interviewees. One of the 10 mentioned the medium increasing their likelihood of clicking on the link. S/he mentioned that it is “not that abnormal for [them] to receive” text messages from their home city" (T13). However, most targets saw the text medium as a red flag, even when they were inclined to believe the content. This was sometimes because of warnings from trainings they had received. One explained that “at work we were warned that all, like, valid messages will be through emails” (T25) and “This one, [employer] is not going to reach out to me via text for security stuff, period” (T08). T37 mentioned that they don’t have a “company texting deal,” making them more suspicious of SMS messages. Another men-tioned that “it’s very rare that people message me, you know, we talk over Slack or we talk over WhatsApp, maybe through email” (T31). Others noted that other communication tools, such as an online learning platform (T09) or a genealogy website (T02), would not have sent them a text message. Interestingly, in all these cases, the technology communications medium was associated with the source of the message. In other words, receiving a text message was not abnormal in and of itself, but receiving a text message from a particular source was deemed either appropriate or not.
5.2.4 Messaging Style
A quarter of respondents (6 of 25) discussed the role of emojis in the context of message style. All six asserted that emojis diminished the credibility of the message. T34 expressed that “I think a lot of emojis is kind of something that usually is a red flag for me.” Multiple respondents conveyed that they did not anticipate receiving emoji-filled messages from senders claiming to represent professional organizations. For instance, T27 said that in the case of a legitimate organization such as a city biking club, emojis would be “out of place." Similarly, T20 explained that if a message claimed to be from their university, the presence of emojis would cause them to feel “thrown off." From the perspective of participants, the use of emojis in messages, particularly those seemingly sent by professional groups, caused such messages to lose legitimacy.
@@ 5.2.5 Scarcity Principle
The remaining participants (6 of 25) said they were less likely to click on the link if it included tactics such as fear and urgency. These participants explained that time sensitive or urgent messages feel like a warning sign. “That’s always a way of like, okay, they’re trying to get you to shut off your logical brain. Put yourself into ’Oh, my gosh, we have to do this right now.’ Yeah. Right.” Another participant confirmed, “If it seems urgent to you, you’re not going to click it because it’s very dangerous” (T13). Similarly, T41 explained, “I don’t like anything where it’s like hurry, offer and send because that always makes me wary” (T07). Similar to the their concerns with time-sensitive messages, participants also discussed how fear inciting messages can cause them to logically analyze the message at a deeper level. One explained, “To the message itself, I was a little bit unsure about just like, the fact that it was so time sensitive, and that it was like, kind of threatening to lock me out of all of my educational accounts if I didn’t do what it asked for. Because that’s not typically what I’ve seen be the case” (T16). And another stated, “Yeah. And it’s the idea of a new virus for young dogs. I’m going that’s fear mongering” (T01). The scarcity principle has long been used in literature to deceive readers. For the spear phishing messages in this study, participants had mixed reactions. Some felt the intended pressure to click, while others saw the language as manipulative and as a warning sign marking potential danger.
5.2.6 Context Inaccuracies
The presence of inaccuracies in the messages was mentioned by 28% (7 of 25) of the targets. These inaccuracies contributed to perceptions about message credibility. Participants identified discrepancies related to the names or characteristics of individuals or enti-ties within their professional spheres. For example, two targets remarked about the message mentioning a colleague that does not exist, stating “there’s no Mike at work” (T37) or “there’s no one named Sarah on the [my university] instructional design team” (T18). Another participant expressed confusion about being questioned on matters not relevant to their professional duties, saying that they “don’t deal with payables. So why is that asking me regarding payment?” (T25). Another target noted discrepencies related to job responsibilities, asserting that such library-related decisions would “go through me at the library”, making any logistical deviation “immediately sus-picious” (T01). Regarding messages related to their hobbies, one participant raised skepticism about the message claiming a “great deal for a hiking trail” which the participant knows through experience doesn’t actually require payment for use (T30). Another target was also suspicious of the message being from a library that they “just don’t use… so they would not know me to contact me” (T08). In summary, the instances where inaccuracies were mentioned ranged from misrepresented personnel and responsibilities to factual discrepancies about the targets’ affiliations, activities, and expectations.